OpenStudy (anonymous):

find the square root of25/49

OpenStudy (tfht):

\[\sqrt{25}/\sqrt{49}\]

OpenStudy (tfht):

simplify that

OpenStudy (the_fizicx99):

\[\sqrt{\frac{ 25 }{ 49 }} = \sqrt{25} \div \sqrt{49}\] What is \[\sqrt{25}\] ?